# An Ontology-Driven Graph RAG for Legal Norms: A Hierarchical, Temporal, and Deterministic Approach

### v4

@@ Abstract
Retrieval-Augmented Generation (RAG) systems in the legal domain face a critical challenge: standard, flat-text retrieval is blind to the hierarchical, diachronic, and causal structure of law, leading to anachronistic and unreliable answers. This paper introduces an ontology-driven Graph RAG framework designed to overcome these limitations. We ground our knowledge graph in a formal, LRMoo-inspired model that distinguishes abstract legal Works from their versioned Expressions. We model temporal states as efficient aggregations that reuse the versioned expressions (CTVs) of unchanged components, and we reify legislative events as first-class Action nodes to make causality explicit and queryable. This structured backbone enables a unified, planner-guided query strategy that applies explicit policies to deterministically resolve complex requests for (i) point-in-time retrieval, (ii) hierarchical impact analysis, and (iii) auditable provenance reconstruction. Through a case study on the Brazilian Constitution, we demonstrate how this approach provides a verifiable, temporally-correct substrate for LLMs, enabling higher-order analytical capabilities while drastically reducing the risk of factual errors. The result is a practical framework for building more trustworthy and explainable legal AI systems.
@@ 1 Introduction
  • A multi-layered representation of abstract legal Works and their versioned, time-stamped Expressions (Temporal and Language Versions).
  • An efficient aggregation model for propagating changes hierarchically without data redundancy.
  • The reification of legislative events as queryable Action nodes, making causality a first-class citizen in the graph.
This structured approach enables a set of generic and deterministic query patterns capable of performing complex temporal and hierarchical analyses that are infeasible for standard RAG systems.
@@ 2 Related Work
Our work introduces Graph RAG for Legal Norms, an architecture that addresses these gaps by: (a) grounding the graph in a formal, LRMoo-inspired ontology that explicitly represents the hierarchy and versioning of legal texts; (b) enabling deterministic, temporally-aware retrieval with auditable provenance (e.g., “what was the applicable text on 2018-05-01?”); and (c) supporting both local (article-level) and global (corpus-level) analysis through structure-aware aggregation. This proposal combines principles from Graph RAG with temporal KG modeling and normative engineering standards. In doing so, we aim to directly address gaps (i)–(iii) above, while also creating a set of evaluation tasks that test temporal retrieval and provenance accuracy.
Finally, recent work proposing hybrid pipelines (KG + RAG + vectors) in the legal domain is highly inspirational for our experimental design. It validates the hypothesis that fusing structured and textual representations reduces hallucinations and improves explainability—but does not yet, in an integrated manner, address the granularity and temporality requirements that legislation imposes [14].
@@
In sum, while a solid foundation of RAG and TKG techniques exists, along with initial applications to the legal domain, a systematic integration of (1) hierarchical and versioned representation of norms, (2) temporally-aware retrieval with provenance, and (3) scalable Graph-RAG pipelines for both local and global questions is still lacking—gaps that the approach proposed in this paper aims to fill.
@@ 3 The Proposed Framework
To address the challenges of representing and retrieving legal knowledge, we propose an ontology-driven framework that adapts and extends the Graph RAG paradigm. Our approach is designed to deterministically model the hierarchical, temporal, and linguistic complexities of legal norms. This section details the framework, starting from its formal ontological foundation, moving to the graph construction process, and finally elaborating on the key adaptations that enable precise, structure-aware, and temporally-correct retrieval.
3.1 Ontological Foundation: A Multi-Layered Model
At the core of our framework lies a formal model grounded in the IFLA LRMoo ontology [4], an object-oriented representation of the IFLA Library Reference Model (LRM). Our choice of LRMoo over other standards like Akoma Ntoso [11] is deliberate. While Akoma Ntoso provides a powerful XML schema for encoding the structure of legal documents, it represents the conceptual layers of the FRBR/LRM model (Work, Expression, Manifestation) primarily as metadata identifiers within the document’s ‘<meta>‘ block. Our framework, in contrast, elevates these concepts from metadata tags to first-class ontological entities. This direct, structural representation of the Work/Expression distinction is fundamental to our granular temporal modeling approach, as it allows us to build explicit, traversable chains of versions within the graph itself, rather than relying solely on the interpretation of metadata.
Grounding our graph structure in the formal model established in [2], we define four primary entity types that are instantiated as nodes:
@@ 3.2 Graph Construction: From Text to a Structured Knowledge Graph
This creates a clear separation of concerns. The CTV represents what the component said at a point in time, while the CLV represents how it was said in a particular language. Consequently, the text chunk/segment itself—which we term a Text Unit—is logically and exclusively associated with the most specific layer: the Language Version node. This design ensures that every piece of retrievable content is unambiguously tied to both a semantic state (the CTV) and a linguistic expression (the CLV), as depicted in Figure 3. This process results in a knowledge graph that provides a verifiable, point-in-time "ground truth" of the law.
@@ 3.3 Temporal Versions as Aggregations, Not Compositions
Figure 6: Diagram of aggregation relationships (orange) between Temporal Versions, illustrating how child CTVs can be reused by multiple parent CTVs at different times.
@@ 4 Case Study: Modeling and Querying Legal Evolution
4.1 Dataset and Implementation
The case study uses the official texts of the Brazilian Federal Constitution (1988) and all subsequent Constitutional Amendments enacted as of early 2024. Documents were obtained from official repositories and preprocessed into structure-aware segments (titles, articles, paragraphs, items) as input to the Graph RAG pipeline. Our implementation for this study integrates three core components:
@@ 4.1 Dataset and Implementation
    A Relational-Vector Knowledge Store: For this study, our knowledge base was implemented in Oracle Database 23ai. The graph structure of our ontological model was realized over the relational model: Theme, Norm_Component, Temporal Version, Language Version, and Action entities were stored as tables, and their relationships (e.g., hierarchy, versioning) were represented by auxiliary tables and foreign key constraints. The Text Unit table included a dedicated vector column to store the text embedding alongside its content. This architecture demonstrates that the proposed graph model can be effectively implemented on modern relational-vector databases, without requiring a dedicated property graph engine.
@@
    An Embedding Model: The vector representations for all Text Units were generated using the Qwen3 Embedding model (256 dimensions). As a powerful open-source model, it was chosen not only for its strong performance on multilingual text comprehension tasks but also because it can be self-hosted, ensuring data privacy and control within an organization’s own infrastructure.
@@
    An LLM-based Generation Module: The final synthesis of natural language responses, based on the retrieved context, was performed by Google’s Gemini 2.5 Flash, selected for its long-context capabilities and efficiency.
While our implementation leverages the efficiency of a converged database, the framework’s architecture is tool-agnostic. It could be readily implemented with other combinations of dedicated graph databases (e.g., Neo4j), vector stores (e.g., Pinecone, FAISS), and LLMs.
The knowledge graph was populated through a semi-automated pipeline. We processed the official, multi-temporal compiled version of the Constitution (and subsequent Constitutional Amendments), applying the semantic segmentation rules outlined in Section 3.2. The legislative events described in each Constitutional Amendment were then modeled as Action nodes, linking the source provisions to the target components they modified. This process resulted in a fine-grained, diachronic representation of the entire legislative history of the Constitution.
@@ 4.2 Modeling a Legislative Amendment: The Case of Article 6
To illustrate our model in action, we focus on the caput of Article 6 of the Constitution, which enumerates social rights. This provision is an ideal test case, having been amended multiple times. We will analyze its state immediately before and after the enactment of Constitutional Amendment (CA) No. 26 of February 14, 2000, which added the right to "housing" (moradia) to the list, as illustred in Figure 11.
@@
Following our framework, this legislative event is modeled as follows:
@@
As a result, the graph now contains a deterministic and fully traceable record of this specific legislative change. The cause (the command from CA No. 26, represented by the Action node) is explicitly linked to its precise effect: the termination of one textual state and the creation of another. This fine-grained, event-centric representation, previously illustrated in Figure 7, is the key to enabling the accurate temporal queries we demonstrate next.
@@ 4.3 Qualitative Evaluation: Answering Complex Temporal Queries
The ultimate value of our framework lies in its ability to answer complex legal queries that are contingent on the law’s precise state at a specific point in time. We evaluate this capability qualitatively by contrasting the behavior of our ontology-driven system with that of a "standard RAG" baseline. For this case study, a standard RAG is defined as a system operating on a flat index of text chunks derived from the current version of the Constitution, lacking any explicit model of the document’s diachronic structure.
@@
Without temporal awareness, a standard RAG system retrieves the current text most semantically similar to the query. Even when the vector index also contains the contemporary texts of amending acts that describe changes to Article 6, retrieval remains unconstrained by legal validity intervals; the LLM therefore tends to conflate contemporary descriptions with the target historical date. In practice, it surfaces modern content and produces anachronistic results (e.g., including rights such as housing ("moradia") or food ("alimentação" that were introduced later). This illustrates a known limitation of flat retrieval systems that lack an explicit, versioned representation of diachronic change.
Ontology-Driven Graph RAG (generic, deterministic)
Our ontology-driven Graph RAG framework resolves this class of queries through a deterministic multi-step plan, which generalizes to any legal provision and any temporal scope:
@@
    • Temporal scope: interprets “in 1999” as a validity constraint for the date $t$ (e.g., any $t\in[1999\text{-}01\text{-}01,1999\text{-}12\text{-}31]$). This constraint is formalized as an interval condition, avoiding reliance on arbitrary cutoffs such as 31 December. If the query lacks an explicit temporal constraint (e.g., asking "what are the rights…"), the planner defaults the temporal scope to the current system date (‘t = now()‘), effectively retrieving the most up-to-date version.
@@
    Temporal scoping via graph traversal. Starting from the ComponentID, the system traverses its chain of Component Temporal Versions (CTVs). It deterministically selects the version whose validity interval satisfies:
@@
    To ensure determinism when multiple CTVs exist within the year, the system applies a temporal-resolution policy (default $\operatorname{SnapshotLast}$): it chooses $t^{\star}=\sup\{t\in[t_{1},t_{2}]\}$ and returns the CTV valid at $t^{\star}$. The policy is disclosed with the answer. This step is both efficient and unambiguous thanks to the aggregation model (Section 3.3), where temporal versions of unchanged child components are reused rather than duplicated, ensuring a single, canonical representation for any component at any point in time.
@@
    Retrieval of the historically valid text units. From the identified CTV, the system retrieves the corresponding Component Language Version (CLV) in the requested language, and selects the associated Text Units that are relevant to the query scope (e.g., caput and enumerated items if the query asks for “listed rights”).
@@
This strategy naturally supports: (i) temporal intervals ($t\in[t_{1},t_{2}]$), (ii) distinctions between promulgation and validity dates, (iii) multilingual retrieval with fallback policies, and (iv) broader or narrower scopes (e.g., Chapters, Titles) simply by changing the entry node and traversal depth. This includes handling queries about the current state of the law, where the temporal scope simply defaults to the present date. The algorithm remains unchanged, ensuring scalability.
@@
Ontology-Driven Graph RAG (structure-aware)
Our framework handles such structural-temporal queries by treating the legal hierarchy as a primary queryable structure. The process is deterministic:
@@
    Scope Identification: A query planner first identifies the structural target (the Component node for "Chapter II") and the temporal window ("after 2010"). It also applies a default policy for resolving scope membership—typically a "snapshot-anchored" approach, which considers all components that belonged to Chapter II at the beginning of the time window. It also applies a scope-membership policy—by default, a snapshot-anchored cohort at the start of the window. Alternative policies (action-time, lifetime) can be selected when reorganizations or cumulative audits are desired; the chosen policy is stated alongside the result.
@@
    Descendant Component Retrieval: The system traverses the graph downwards from the "Chapter II" node to gather the set of all its descendant components (e.g., Art. 6, Art. 7, etc.) that satisfy the membership policy.
@@
    Causal Event Aggregation: For this set of components, the system retrieves all Action nodes that affected them within the specified time window.
@@
User Query (example): "Trace the full legislative lineage that introduced the term ‘food’ in the caput of Article 6 of the Brazilian Constitution."
@@
Ontology-Driven Graph RAG (strategy-aware)
Our framework’s query planner selects the most efficient execution strategy based on the query’s constraints. As this query provides both a structural ("Article 6") and a textual ("food") target, the planner opts for a structure-first approach to dramatically narrow the search space:
@@
    Query Analysis and Constraint Extraction: The planner first parses the user’s query into a set of structured constraints, without committing to an execution order. This includes identifying and canonicalizing the structural target (‘ComponentID‘ for "Art. 6 caput") and the textual target ("food").
@@
    Hierarchical Scope Resolution: The system takes the identified structural target (‘ComponentID‘ for "Art. 6 caput") and resolves the full search scope. This scope includes the component itself and the entire descendant subtree beneath it in the legal hierarchy. This ensures that if a user specifies a high-level component (e.g., "Article 6"), the search will correctly include all its children (caput, paragraphs, items, etc.).
@@
    • Effect: Inserted the term "food". Textual Diff: + "food".

### v5

# An Ontology-Driven Graph RAG for Legal Norms: A Structural, Temporal, and Deterministic Approach

@@ Abstract
Retrieval-Augmented Generation (RAG) systems in the legal domain face a critical challenge: standard, flat-text retrieval is blind to the hierarchical, diachronic, and causal structure of law, leading to anachronistic and unreliable answers. This paper introduces the Structure-Aware Temporal Graph RAG (SAT-Graph RAG), an ontology-driven framework designed to overcome these limitations by explicitly modeling the formal structure and diachronic nature of legal norms. We ground our knowledge graph in a formal, LRMoo-inspired model that distinguishes abstract legal Works from their versioned Expressions. We model temporal states as efficient aggregations that reuse the versioned expressions (CTVs) of unchanged components, and we reify legislative events as first-class Action nodes to make causality explicit and queryable. This structured backbone enables a unified, planner-guided query strategy that applies explicit policies to deterministically resolve complex requests for (i) point-in-time retrieval, (ii) hierarchical impact analysis, and (iii) auditable provenance reconstruction. Through a case study on the Brazilian Constitution, we demonstrate how this approach provides a verifiable, temporally-correct substrate for LLMs, enabling higher-order analytical capabilities while drastically reducing the risk of factual errors. The result is a practical framework for building more trustworthy and explainable legal AI systems.
@@ 1 Introduction
  • A formal, multi-layered representation, building upon the model in [2], that separates abstract legal Works from their versioned, time-stamped Expressions.
  • An efficient aggregation model for propagating changes hierarchically, which avoids data redundancy by reusing unchanged versioned components.
  • The reification of legislative events as first-class, retrievable entities, making causality a directly queryable aspect of the law’s history.
  • A multi-aspect retrieval approach that textualizes structured metadata and events, making the context of a norm as searchable as its content.
This structured approach enables deterministic query patterns for complex temporal and analyses that are infeasible for standard RAG systems. To better articulate these contributions, we find it useful to conceptually distinguish two core facets of our framework. We refer to the use of the intrinsic document hierarchy—its titles, chapters, and articles—as the graph’s backbone as the Structure-Aware Graph RAG perspective. This contrasts with standard Graph RAG, which derives a semantic hierarchy from content rather than adopting the document’s formal schema. Complementing our approach, we refer to the modeling of the diachronic evolution of these structural components as the Temporal Graph RAG perspective. It is the seamless integration of these two perspectives, grounded in a formal ontology, that forms the basis of our unified Structure-Aware Temporal Graph RAG (SAT-Graph RAG) framework.
@@ 2 Related Work
Our work introduces Graph RAG for Legal Norms, an architecture that addresses these gaps by: (a) grounding the graph in a formal, LRMoo-inspired ontology that explicitly represents the hierarchy and versioning of legal texts; (b) enabling deterministic, temporally-aware retrieval with auditable provenance (e.g., “what was the applicable text on 2018-05-01?”); and (c) supporting both local (article-level) and global (corpus-level) analysis through structure-aware aggregation. This proposal combines principles from Graph RAG with temporal KG modeling and normative engineering standards. In doing so, we aim to directly address gaps (i) and (ii), and lay the groundwork for addressing gap (iii) by creating a set of evaluation tasks that test temporal retrieval and provenance accuracy.
Finally, our framework aligns with a growing body of research proposing hybrid pipelines (KG + RAG + vectors) in the legal domain. This line of work validates the hypothesis that fusing structured and textual representations reduces hallucinations and improves explainability. Our approach can be seen as a sophisticated instantiation of this principle, but one that is specifically designed to address the deep granularity and temporality requirements that legislation imposes, which current hybrid models do not yet fully address [14].
@@
In sum, while a solid foundation of RAG and TKG techniques exists, along with initial applications to the legal domain, a systematic integration of (1) a representation of the norm’s formal structure and its versioning, (2) temporally-aware retrieval with provenance, and (3) scalable Graph-RAG pipelines for both local and global questions is still lacking—gaps that the approach proposed in this paper aims to fill.
@@ 3 The Proposed Framework
Our framework models the formal-structural, temporal, and linguistic complexities of legal norms through an ontology-driven graph structure. Its design is motivated by a logical progression that addresses the shortcomings of existing Graph RAG approaches when applied to structured, diachronic domains like legal norms.
3.1 Design Rationale: From Entity-Centric to Structural-Temporal RAG
Standard Graph RAG approaches primarily focus on an entity-centric model [3]. In this model, the graph is constructed by extracting named entities (e.g., people, organizations) and key concepts from the text, which become nodes, while their co-occurrences or inferred relationships become edges. While powerful for discovering thematic connections within content, this approach is fundamentally blind to the formal, hierarchical architecture of the document itself. For legal texts, where the position of a provision within a chapter or title defines its scope and meaning, this structural blindness is a critical limitation.
Beyond its structural blindness, the standard entity-centric model faces a more fundamental challenge in the statutory domain. Named Entity Recognition (NER) systems are optimized to extract proper nouns (e.g., persons, organizations, locations), which are notably sparse in the abstract and impersonal text of legal norms. The true semantic fabric of law is woven from legal concepts (e.g., "administrative act," "due process"), which are not captured by standard NER models. A direct application of an entity-extraction pipeline would therefore yield a sparse and semantically impoverished knowledge graph. This further justifies our decision to pivot from a content-driven, bottom-up approach to a top-down model that leverages the document’s formal structure—an ever-present and high-value feature—as the graph’s primary backbone.
Then, the first key design decision of our framework is to adopt a structural perspective. Instead of content-derived entities, we model the intrinsic structure of the legal norm as the graph’s backbone. Hierarchical components of the document—such as Titles, Chapters, Articles, and Items—are instantiated as nodes, and their parent-child relationships form the primary edges. This shift allows queries to target not just the text’s content, but the architecture of the knowledge itself.
However, a purely structural graph would still be temporally-naïve, representing only a single state of the law and failing to capture its diachronic nature. The final and most critical layer of our design is therefore the introduction of the temporal perspective. We elevate this structural model by ensuring that each version of a structural component over time becomes a distinct, addressable node in the graph. As detailed below, this is achieved by distinguishing abstract components from their time-stamped Temporal Versions (CTVs). It is the synthesis of these structural and temporal dimensions that forms our unified Structure-Aware Temporal Graph RAG (SAT-Graph RAG) model, enabling a retrieval mechanism with the precision and determinism required for the legal domain.
3.2 Ontological Foundation and Graph Construction
At the core of our framework lies a formal model grounded in the IFLA LRMoo ontology [4], an object-oriented representation of the IFLA Library Reference Model (LRM). Our choice of LRMoo over other standards like Akoma Ntoso [11] is deliberate. While Akoma Ntoso provides a powerful XML schema for encoding the structure of legal documents, it represents the conceptual layers of the FRBR/LRM model (Work, Expression, Manifestation) primarily as metadata identifiers within the document’s ‘<meta>‘ block.
Our framework, in contrast, elevates these concepts from metadata tags to first-class ontological entities. We therefore ground our graph structure in the formal model established in [2]. This model’s direct, structural representation of the Work/Expression distinction is fundamental to our granular temporal modeling approach, as it allows us to build explicit, traversable chains of versions within the graph itself, rather than relying solely on the interpretation of metadata. Following this model, we define four primary entity types that are instantiated as nodes:
@@ 3.3 Graph Construction: From Text to a Structured Knowledge Graph
This creates a clear separation of concerns. The CTV represents the conceptual substance of the component at a point in time, while the CLV represents the specific wording used to express it in a particular language. Consequently, the text chunk/segment itself—which we term a Text Unit—is logically and exclusively associated with the most specific layer: the Language Version node. This design ensures that every piece of retrievable content is unambiguously tied to both a semantic state (the CTV) and a linguistic expression (the CLV), as depicted in Figure 3. This process results in a knowledge graph that provides a verifiable, point-in-time "ground truth" of the law.
@@ 3.4 Temporal Versions as Aggregations, Not Compositions
Figure 6: Diagram of aggregation relationships (orange) between Temporal Versions, illustrating how child CTVs are reused by multiple parent CTVs at different times. For instance, the ‘v2‘ Temporal Version of ‘c1‘ aggregates the new ‘v2‘ CTV from Component ‘C2‘ but reuses the pre-existing ‘v1‘ CTV from the unchanged Component ‘C3‘.
@@ 4 Case Study: Modeling and Querying Legal Evolution
4.1 Dataset and Conceptual Architecture
The case study uses the official texts of the Brazilian Federal Constitution (1988) and all subsequent Constitutional Amendments enacted as of early 2024. The documents would serve as input for a pipeline that preprocesses them into structure-aware segments (titles, articles, paragraphs, items). A potential implementation of our framework would integrate three core components:
@@ 4.1 Dataset and Conceptual Architecture
    A Relational-Vector Knowledge Store: For this study, our knowledge base could be implemented in a modern relational-vector database such as Oracle Database 23ai. The graph structure of our ontological model could be realized over the relational model: Theme, Norm_Component, Temporal Version, Language Version, and Action entities would be stored as tables, and their relationships (e.g., hierarchy, versioning) represented by auxiliary tables and foreign key constraints. The Text Unit table would include a dedicated vector column to store the text embedding alongside its content. This architecture demonstrates that the proposed graph model can be effectively implemented on such databases, without requiring a dedicated property graph engine.
@@
    An Embedding Model: The vector representations for all Text Units could be generated using a powerful open-source model such as Qwen3 Embedding. Such a model is a strong candidate not only for its performance on multilingual text comprehension but also because it can be self-hosted, ensuring data privacy and control.‘
@@
    An LLM-based Generation Module: The final synthesis of natural language responses, based on the retrieved context, could be performed by a model like Google’s Gemini 1.5 Flash, selected for its long-context capabilities and efficiency.
While this conceptual architecture could leverage the efficiency of a converged database, the framework itself is tool-agnostic. It could be readily implemented with other combinations of dedicated graph databases (e.g., Neo4j), vector stores (e.g., Pinecone, FAISS), and LLMs.
The knowledge graph would be populated through a semi-automated pipeline. This process would involve processing the official, multi-temporal compiled version of the Constitution (and subsequent Constitutional Amendments), applying the semantic segmentation rules outlined in Section 3.2. The legislative events described in each Constitutional Amendment would then be modeled as Action nodes, linking the source provisions to the target components they modified. This process would result in a fine-grained, diachronic representation of the entire legislative history of the Constitution.
@@ 4.2 Modeling a Legislative Amendment: The Case of Article 6
To illustrate the application of our model, we focus on the caput of Article 6 of the Constitution, which enumerates social rights. This provision is an ideal test case, having been amended multiple times, as showed in Figure 11.
@@
We will analyze its state immediately before and after the enactment of Constitutional Amendment (CA) No. 26 of February 14, 2000, which added the right to "housing" (a moradia) to the list. Following our framework, this legislative event is modeled as follows:
@@
As a result, the graph would contain a deterministic and fully traceable record of this specific legislative change. The cause (the command from CA No. 26, represented by the Action node) is explicitly linked to its precise effect: the termination of one textual state and the creation of another. This fine-grained, event-centric representation, previously illustrated in Figure 7, is the key to enabling the accurate temporal queries we demonstrate next.
@@ 4.3 Qualitative Evaluation: Answering Complex Temporal Queries
The ultimate value of our framework lies in its ability to answer complex legal queries that are contingent on the law’s precise state at a specific point in time. We evaluate this capability qualitatively by contrasting the designed behavior of our ontology-driven system with that of a "standard RAG" baseline. For this case study, a standard RAG is defined as a system operating on a flat index of text chunks derived from the current version of the Constitution, lacking any explicit model of the document’s diachronic structure.
@@
Without temporal awareness, a standard RAG system would retrieve the current text most semantically similar to the query. Even when the vector index also contains the contemporary texts of amending acts that describe changes to Article 6, retrieval remains unconstrained by legal validity intervals; the LLM therefore would tend to conflate contemporary descriptions with the target historical date. In practice, it would surface modern content and produce anachronistic results (e.g., including rights such as housing ("a moradia") or food ("a alimentação") that were introduced later. This illustrates a known limitation of flat retrieval systems that lack an explicit, versioned representation of diachronic change.
Ontology-Driven SAT Graph RAG (generic, deterministic)
Our ontology-driven SAT Graph RAG framework is designed to resolve this class of queries through a deterministic multi-step plan, which generalizes to any legal provision and any temporal scope:
@@
    • Temporal scope: interprets “in 1999” as a validity constraint for the date $t$ (e.g., any $t\in[1999\text{-}01\text{-}01,1999\text{-}12\text{-}31]$). This constraint is formalized as an interval condition, avoiding reliance on arbitrary cutoffs such as 31 December. If the query lacks an explicit temporal constraint (e.g., asking "what are the rights…"), the planner defaults the temporal scope to the current system date (t = now()), effectively retrieving the most up-to-date version.
@@
    Temporal scoping via graph traversal. This is the core deterministic step. Starting from the identified ComponentID ("Art. 6"), the system first resolves the full hierarchical scope by traversing downwards to gather the entry component itself and all its text-bearing descendants (e.g., the nodes for ‘Art. 6‘, its ‘caput‘, and any ‘items‘). With this set of target components defined, the system traverses their chain of Component Temporal Versions (CTVs) to select the versions whose validity interval satisfies:
@@
    This step yields the precise set of CTVs that represent the complete state of Article 6 and its parts in 1999. To ensure determinism when multiple CTVs exist within the year, the system applies a temporal-resolution policy (default $\operatorname{SnapshotLast}$): it chooses $t^{\star}=\sup\{t\in[t_{1},t_{2}]\}$ and returns the CTV valid at $t^{\star}$. The policy is disclosed with the answer. This step is both efficient and unambiguous thanks to the aggregation model (Section 3.4), where temporal versions of unchanged child components are reused rather than duplicated, ensuring a single, canonical representation for any component at any point in time.
@@
    Retrieval of the historically valid text units. From the set of valid CTVs identified in the previous step, the system retrieves the corresponding Component Language Versions (CLVs) in the requested language, and selects the associated Text Units that are relevant to the query scope.
@@
This strategy naturally supports: (i) temporal intervals ($t\in[t_{1},t_{2}]$), (ii) distinctions between promulgation and validity dates, (iii) multilingual retrieval with fallback policies, and (iv) querying diverse scopes simply by changing the entry node. This includes scopes defined by the document’s internal hierarchy (e.g., from a specific item to entire Chapters or Titles) as well as by external, cross-document thematic classifications (e.g., starting from a ’Social Security’ Theme node). The algorithm remains unchanged, handling even queries about the current state of the law by defaulting the temporal scope to the present date, thus ensuring scalability.
@@
Ontology-Driven SAT Graph RAG (structure-aware)
Our framework is designed to handle such structural-temporal queries by treating the legal hierarchy as a primary queryable structure. The process is deterministic:
@@
    Scope Identification and Policy Selection: The query planner first identifies the structural target (the Component node for "Chapter II") and the temporal window ("after 2010"). It also selects a scope-membership policy to handle potential structural reorganizations over time. The default is a "snapshot-anchored" approach, which defines the scope as all components that belonged to Chapter II at the beginning of the time window. Alternative policies (e.g., action-time, lifetime) can be used for different analytical needs, and the chosen policy is always disclosed with the result.
@@
    Descendant Component Retrieval: With the entry node and membership policy defined, the system executes the retrieval. It traverses the graph downwards from the "Chapter II" node, gathering the set of all descendant components (e.g., Art. 6, Art. 7) that satisfy the membership policy selected in the previous step.
@@
    Causal Event Aggregation: For this set of components, the system gathers all legislative events that drove their evolution within the specified time window. This is achieved by retrieving all Action nodes that acted upon the Component Temporal Versions (CTVs) associated with these components.
@@
User Query (example): "Trace the full legislative lineage that introduced the term ’food’ in Article 6 of the Brazilian Constitution."
@@
Ontology-Driven SAT Graph RAG (strategy-aware)
Our framework’s query planner would select the most efficient execution strategy based on the query’s constraints. As this query provides both a structural ("Article 6") and a textual ("food") target, the planner would opt for a structure-first approach to narrow the search space:
@@
    Query Analysis and Constraint Extraction: The planner first parses the user’s query into a set of structured constraints, without committing to an execution order. This includes identifying and canonicalizing the structural target (‘ComponentID‘ for "Art. 6") and the textual target ("food").
@@
    Hierarchical Scope Resolution: The system takes the identified structural target (‘ComponentID‘ for "Art. 6") and resolves the full search scope. This scope includes the component itself and the entire descendant subtree beneath it in the legal hierarchy. This ensures that if a user specifies a high-level component (e.g., "Article 6"), the search will correctly include all its children (caput, paragraphs, items, etc.).
@@
    • Effect: Inserted the term "food".